From equation 4CH3COOH + 12CH3OH <--> 5C4H8O2 + 10H2O,

If 0.5mol of alcohol and 0.2 mol of acid are initially used, what is the equilibrium composition if the equilibrium coefficient(K) is 8. Use the mole balance and then solve it again with the compound balance.
Answer #1

Given: K = 8 I think the reaction is not balanced and the stoichiometric coefficients are wrong. The correct balanced chemica

Here, An = v = -4 -8 + 4 + 8 = 0 (vi is stoichiometric coefficients) For equilibrium reactions, K*p-An = YC4H802* YH2O Үснэсо